OpenStudy (electrochika):

dy/ dx of y=sin^3xtan4x

OpenStudy (anonymous):

you can use the product rule

OpenStudy (amistre64):

good thing it aint the integral

OpenStudy (anonymous):

Product and chain rule.

OpenStudy (anonymous):

@amistre64:shouldn't by parts work then ?

OpenStudy (amistre64):

it might, but i dunno

OpenStudy (electrochika):

I don't know how to multiply it in.

OpenStudy (amistre64):

aint got to multiply it in; product rule is: [fg]' = f'g + fg'

OpenStudy (anonymous):

just find the derivaitve of sin^3x and tan4x seperatly, then you can put them together

OpenStudy (electrochika):

the first times the derivative of the second plus the second times the derivative of the first?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

or if you have more than 2 it just a sum that allows each one to get derived seperately

OpenStudy (amistre64):

[fgh]' = f'gh + fg'h + fgh'

OpenStudy (electrochika):

i set it up in the product rule but now i'm stuck

OpenStudy (amistre64):

what you got?

OpenStudy (anonymous):

\[ 4 \text{Sec}[4 x]^2 \text{Sin}[x]^3+3 \text{Cos}[x] \text{Sin}[x]^2 \text{Tan}[4 x] \] I am feeling lazy :P

OpenStudy (electrochika):

sin^3x((sec^2(4x)) + tan4x(cos^3x) is that even right?

OpenStudy (anonymous):

I gave you the correct answer

OpenStudy (electrochika):

I need to show steps

OpenStudy (anonymous):

OHHHH.. then follow the instructions of amistre...

OpenStudy (anonymous):

what instructions? use \[(fg)'=f'g+g'f\] with \[f(x)=\sin^3(x),f'(x)=3\sin^2(x)\cos(x),g(x)=\tan(4x), g'(x)=4\sec^2(x)\] and put it together